OpenStudy (cindy):

simplify 6(1/2-2/3y) can somebody help me on this problem

OpenStudy (gw2011):

The answer is: 3-4y

OpenStudy (anonymous):

You should probably tell her how you got that.... ;P

OpenStudy (anonymous):

the simplification is obvious..jus mutiply the terms inside the bracket with 6..

OpenStudy (gw2011):

You're probably right that I should show my work. Here it is: (6)(1/2)=3 because 1/2 of 6 is 3 (6)[(-2/3)(y)]=-4y because -2/3 of 6=-4 and you have a y associated with it. Therefore, this problem simplifies to 3-4y. Hopefully, This helps you understand how the answer was arrived at.

OpenStudy (cindy):

but isnt the answer supoose to be a fraction

OpenStudy (gw2011):

No, the answer is not supposed to be a fraction because the fractions actually cancel as you can see from the above work I showed. Do you now understand how this worked?

OpenStudy (cindy):

im still confuse,,sorry

OpenStudy (gw2011):

6*1/2=6/2=3 6*2/3=12/3=4 but 2/3 is a negative so 4 becomes -4 but again there is a y associated with -2/3 so -4 becomes -4y. Does this help?